1/x+1/y=1/z solve for x. I have seen the answer, x=yz/y-z. I need to know how the answer is derived.
Answer by scott8148(6628)

subtracting 1/y ___ 1/x = 1/z - 1/y

combining fractions (common denominator) ___ 1/x = (y / yz) - (z / yz) = (y - z) / yz

taking inverse ___ x = yz / (y - z)